OpenStudy (anonymous):

find the slop, if it exists of the line containing the pair of points (2,8) and (-2,8). I got 3/9 is this correct

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the y values are the same the line is horizontal the slope is zero

OpenStudy (bradely):

slope =y2-y1/x2-x1 =8-8/-2-2 =0

OpenStudy (anonymous):

thank you